OpenStudy (anonymous):

I know this is gonna sound dumb, but, if the area of a square is .81, wouldn't the perimeter be 0.9?

OpenStudy (anonymous):

No, 3.6

OpenStudy (anonymous):

Whaaaa? Thought you squared it?

OpenStudy (cwrw238):

one side would be 0.9 as 0.9^2 = 0.81 perimeter = sum of 4 sides =4 * 0.9 = 3.6

OpenStudy (anonymous):

Side is 0.9, perimeter is 3.6

OpenStudy (anonymous):

Thanks guys. Saved my seventh grade son some time:)